Assume that a company intends to sale product in the market, at a selling price of sh.9 per
unit. The V C is shs.5 per unit and the T F C is sh.2000
Required:
i. Compute the B E P in units and in shs.
ii. Assume that the company intends to make a profit before tax of 20% of sales,
determine the number of units that must be sold.
iii. Assume that the corporate tax rate is 30% and the company has a target profit
of 1640 after tax. Compute the number of units that must be sold to earn this
target profit.
If the company expects to sale 600 units, compute the marginal of safety.
1
Expert's answer
2015-03-05T08:42:58-0500
P = 9, AVC = 5, TFC = 2000
Required:
i. BEP in units is Q = TFC/(P - AVC) = 2000/(9 - 5) = 500 units.
TR = 500*9 = shs 4,500
ii. If the company intends to make a profit before tax of 20% of sales, the number of units that must be sold is:
TP = P*Q - TFC - AVC*Q
Q = TFC/(0.8P - AVC)
Q = 2000/(0.8*9 - 5) = 909.1 units
iii. If the tax rate is 30% and the company has a target profit
of 1640 after tax, the number of units that must be sold to earn this target profit is:
Q = (TP/0.7 + TFC)/(P - AVC) = (1640/0.7 + 2000)/4 = 4342/4 = 1085.7